OpenStudy (andrewkaiser333):

Use Einstein’s energy equation to solve the following problem. Recall that the speed of light is 3 × 108 m/s. If a 5.0-kg stone is completely converted to energy, what is its energy equivalent in joules? 4.5 × 10^17 J 1.5 × 10^10 J 1.5 × 10^9 J 4.5 × 10^15 J

OpenStudy (andrewkaiser333):

@jim_thompson5910

OpenStudy (andrewkaiser333):

@anyone

OpenStudy (anonymous):

\(\sf E= mc^2\) , where m is the mass (5 kg) & c is the speed of light ( 3 × 108 m/s) . All you have to do is substitute the given values.

OpenStudy (andrewkaiser333):

i am not to good with that equation

jimthompson5910 (jim_thompson5910):

Given info \[\Large m = 5\] \[\Large c = 3 \times 10^8\] plug them into the formula \[\Large E = mc^2\] \[\Large E = 5*\left(3 \times 10^8\right)^2\] then evaluate

OpenStudy (andrewkaiser333):

oh ok

OpenStudy (andrewkaiser333):

(15*50^8)^2

jimthompson5910 (jim_thompson5910):

btw, \[\Large 3 \times 10^8 = 300,000,000\] which is the number `300 million`

jimthompson5910 (jim_thompson5910):

PEMDAS says you have to do exponents before multiplication

jimthompson5910 (jim_thompson5910):

you cannot multiply the 5 and 3 because the 3 is under an exponent

OpenStudy (andrewkaiser333):

ok then can u show me? i am confused

OpenStudy (andrewkaiser333):

Like i said not too good with this formula

jimthompson5910 (jim_thompson5910):

\[\Large E = 5*\left(3 \times 10^8\right)^2\] \[\Large E = 5*\left(3\right)^2*\left(10^8\right)^2\] \[\Large E = 45*\left(10^8\right)^2\] I'll let you finish up

OpenStudy (andrewkaiser333):

45*10^10 ?

jimthompson5910 (jim_thompson5910):

you won't add the exponents in this case

OpenStudy (andrewkaiser333):

how?

jimthompson5910 (jim_thompson5910):

when you raise an exponential expression to another exponent, you multiply the exponents

jimthompson5910 (jim_thompson5910):

example: \[\Large (2^3)^5 = 2^{3*5} = 2^{15}\]

OpenStudy (andrewkaiser333):

ok so 8*2= 45*10^16

jimthompson5910 (jim_thompson5910):

now convert \[\Large 45 \times 10^{16}\] to scientific notation

OpenStudy (andrewkaiser333):

4.5*10^16 ???

jimthompson5910 (jim_thompson5910):

no

OpenStudy (andrewkaiser333):

then what??

jimthompson5910 (jim_thompson5910):

it might help to think of 45 as 4.5*10^1

OpenStudy (andrewkaiser333):

ok?

jimthompson5910 (jim_thompson5910):

\[\Large 45*10^{16} = (4.5 \times 10^1)*10^{16} = ??\]

OpenStudy (andrewkaiser333):

the 10^16 is multiplied by 4.5 and 10^1 yes??

jimthompson5910 (jim_thompson5910):

what is 10^1 times 10^16 equal to? in terms of an exponential expression

OpenStudy (andrewkaiser333):

4.5*10^17

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (andrewkaiser333):

ok thanks